# MARFT: Multi-Agent Reinforcement Fine-Tuning

### v3

@@ Abstract
Large Language Model (LLM)-based Multi-Agent Systems (LaMAS) have demonstrated remarkable capabilities in addressing complex, agentic tasks requiring multifaceted reasoning and collaboration, from generating high-quality presentation slides to conducting sophisticated scientific research. Meanwhile, Reinforcement Learning (RL) has been widely recognized for its effectiveness in enhancing agent intelligence, but limited research has investigated the fine-tuning of LaMAS using foundational RL techniques. Moreover, the direct application of conventional Multi-Agent Reinforcement Learning (MARL) methodologies to LaMAS introduces significant challenges, stemming from the unique characteristics and mechanisms inherent to LaMAS. To address these challenges, this article presents a comprehensive study of LLM-based MARL and proposes a novel paradigm termed Multi-Agent Reinforcement Fine-Tuning (MARFT). We introduce a brand-new POMDP called Flex-POMDP, which aligns with the LaMAS optimization in real world applications and a universal algorithmic framework tailored specifically for LaMAS, outlining the conceptual foundations, key distinctions, and practical implementation strategies. We begin by reviewing the evolution from traditional RL to Reinforcement Fine-Tuning (RFT), setting the stage for a parallel analysis in the multi-agent domain. In the context of LaMAS, we elucidate critical differences between classical MARL and MARFT, such as asynchronous agent interactions, profile-aware agent design, and heterogeneous architectural configurations, etc. These differences motivate a transition toward a novel, LaMAS-oriented formulation of RFT. Central to this work is the presentation of a robust and scalable MARFT framework. We detail the core algorithm, emphasizing its modularity and adaptability, and provide a complete, open-source implementation to facilitate adoption and further research. The latter sections of the paper explore real-world application perspectives and opening challenges in MARFT, including dynamic environment modeling, sample inefficiency, and the current lack of cohesive frameworks. By bridging theoretical underpinnings with practical methodologies, this work aims to serve as a roadmap for researchers seeking to advance MARFT toward resilient, adaptive, and human-aligned solutions in agentic systems. Our implementation of the proposed framework is publicly available at: https://github.com/jwliao-ai/MARFT.
@@ 8             $\mathcal{D}\leftarrow\mathcal{D}\cup\{\left(s,a,s^{\prime},r\right)\}$
11      sample a batch =$\mathcal{B}=\{(s,a,s^{\prime},r)\}\sim\mathcal{D}$
@@ 2.2 LLM-based Multi-Agent Systems (LaMAS)
  • Parameter fine-tuning. A more sophisticated approach involves updating agent parameters or network configurations to optimize performance. This paradigm integrates deep learning and specific methodologies, such as multi-agent debating to generate high-quality training datasets (Subramaniam et al., 2025) or fine-tuning programming modules for improved workflow organization (Khattab et al., 2024). CORY (Ma et al., 2024) duplicates the LLM into two agents, i.e., a pioneer and an observer, that interact through knowledge transfer and role exchange, improving training stability and performance. However, many parameter fine-tuning methods lack a theoretical foundation in RL, relying instead on ad-hoc techniques. SiriuS (Zhao et al., 2025), a self-improving, multi-agent optimization framework using bootstrapped reasoning, iteratively refines agent behaviors by leveraging successful reasoning trajectories and augmenting unsuccessful ones. The most related work, MAPoRL (Park et al., 2025), rooted in MARL, pioneers in applying multi-agent PPO to post-train LLMs, but does not consider the unique characteristics of LaMAS and assumes homogeneous agent behavior through collaborative debate, diverging from real-world scenarios where agents often perform diverse or orthogonal roles to solve complex problems. This article will focus on RFT and MARL-related methodologies in subsequent sections, providing a principled paradigm for general LaMAS reinforcement.
@@ 2.3.2 Main Approaches of MARL
(Multi-Agent Advantage Decomposition Theorem (Kuba et al., 2021)). For any predefined permutation of $n$ agents, for any joint observation $\bm{o}\in\bm{\mathcal{O}}$ and joint action $\bm{a}\in\bm{\mathcal{A}}$, the following always holds:
$A^{\bm{\pi}}(\bm{o},\bm{a}^{1:n})=\sum_{m=1}^{n}A^{\bm{\pi}}(\bm{o},\bm{a}^{1:m-1},\bm{a}^{m}).$ (17)
@@
$L(\theta)=\frac{1}{nT}\sum_{i=1}^{n}\sum_{t=0}^{T-1}\min\left[\bm{r}_{t}^{i}(\theta)\hat{A}_{t},\text{clip}\left(\bm{r}_{t}^{i}(\theta),1\pm\epsilon\right)\hat{A}_{t}\right],\text{where}\ \bm{r}_{t}^{i}(\theta)=\frac{\pi_{\theta}^{i}\left(a_{t}^{i}|\hat{\bm{o}}_{t}^{1:n},\hat{\bm{a}}_{t}^{1:i-1}\right)}{\pi_{\theta_{\text{old}}}^{i}\left(a_{t}^{i}|\hat{\bm{o}}_{t}^{1:n},\hat{\bm{a}}_{t}^{1:i-1}\right)}.$ (19)
@@ 3 MARFT Overview
3.1 Problem Statement and Flex-POMDP
@@
Figure 2: A detailed illustration of the dynamics of a Flex-POMDP. The dependency function (dashed purple line) can vary across timesteps.
To enhance problem formulation and accommodate MARFT, we propose Flexible Partially Observable Markov Decision Process (Flex-POMDP), illustrated in Figure 2 and denoted as $\langle\mathcal{V},\mathcal{N},{\mathcal{S}},\bm{\mathcal{O}},\bm{\mathcal{A}},\mathcal{T},\mathcal{R},\gamma,\mathcal{D}\rangle$, where $\mathcal{N}=\{1,...,n\}$ is the agent composition of a LaMAS with some organization, i.e., some order constraint among agents. $\mathcal{V},\mathcal{S},\bm{\mathcal{O}},\bm{\mathcal{A}},\mathcal{T},\mathcal{R}\ \text{and}\ \gamma$ can be borrowed from Section 2.3.1. Flex-POMDP extends the DEC-POMDP (Oliehoek & Amato, 2016) by introducing a dependency function $\mathcal{D}\colon\bm{\mathcal{A}}\times\bm{\mathcal{A}}\rightarrow\{0,1\}$, which explicitly models inter-agent dependencies at the level of action execution or outcome. Specifically, $\mathcal{D}(a^{i},a^{j})=1$ indicates that the action of agent $j$ is conditionally dependent on the action of agent $i$. As a result, the decision-making process of agent $j$ is influenced not only by its own observation but also by the actions of all agents $k$ such that $k\in\{l|\mathcal{D}(a^{l},a^{j})=1\}$. These dependencies can be integrated into agent $j$’s input via concatenation or other fusion methods, thereby forming an enriched "rollout observation". Crucially, the dependency function $\mathcal{D}$ can vary across timesteps, potentially governed by an orchestration agent or a dynamic routing mechanism. This flexibility captures the evolving coordination structures in LaMAS and embodies the "flexible" nature of the proposed framework. Notably, when $\mathcal{D}(a^{i},a^{j})=0$ for all $i,j$, Flex-POMDP reduces to the standard DEC-POMDP, thereby demonstrating the generality of the formulation.
@@ 3.3 Differences and Resulting Challenges
Action Asynchronicity Synchronous Asynchronous
Ultilities Task-specific only Agentic with original language ability
@@
Heterogeneity Parameters In-out formats, external databases, params
System Organization Static Highly dynamic
@@
Action Asynchronicity
One of the most significant differences between conventional MARL and MARFT is the nature of agent actions. In traditional cooperative MARL, agents typically act simultaneously. However, in MARFT, agents often act asynchronously. In some cases, the actions of certain agents may even depend on the outcomes of other agents’ actions. For example, in a collaborative coding assistant system, one LLM agent generates a function prototype, and another agent asynchronously refines it based on the first agent’s output, demonstrating both asynchronicity and result dependency. As a result, conventional MARL methods, which assume synchronous actions, may not be directly applicable or effective in LLM-based multi-agent systems.
@@ 4 Multi-Agent Reinforcement Fine-Tuning
4.1 MARFT at Different Granularity
4.1.1 Action-level Fine-Tuning
Action-level MARFT
Inheriting the SM-style re-modeling via the multi-agent advantage decomposition theorem introduced in the last section, a naive variant of MARFT, i.e., Action-level MARFT, is developed and unifies the dynamic LaMAS decision processes perfectly. Its procedure is demonstrated in Figure 3.
Figure 3: The procedure of Multi-Agent Reinforcement Fine-Tuning. Inference and training are conducted in an alternating manner. The dynamic organization of agents in the LLM-based MAS depends on the dependency function $\mathcal{D}$. As elaborated in Section 3.1, this function is managed by an orchestration agent or a routing mechanism. Each agent within the MAS can have its own private APIs, tool pool, database, and other resources.
@@ 4.1.1 Action-level Fine-Tuning
Given any organizational or task-solving workflow of a LaMAS, theorem (17) allows MARFT to reframe it as a sequential decision-making process. When collecting interactive trajectories, the LaMAS generates actions with an auto-regressive problem-solving style. The "encoder" constructs local roll-out observations using agent profiles and other relevant information, while the "decoder" combines these observations with dependent predecessors’ actions to generate the current agent’s action $a^{i}\sim p_{\pi^{i}}(a^{i}|\hat{o}^{i},a^{\mathcal{N}(i)})$, where $\mathcal{N}(i)=\left\{j\neq i\mid\mathcal{D}(a^{j},a^{i})=1\right\}$. Then, during training, MARFT adopts an optimization objective in (20). Notably, in naive MARFT, each agent optimizes within a trust region conditioned on predecessor actions, akin to MAT, thus ensuring monotonic improvement as in HAPPO’s sequential update scheme.
$L(\bm{\theta})=\frac{1}{nT}\sum_{i=1}^{n}\sum_{t=0}^{T-1}\min\left[\bm{r}_{t}^{i}(\bm{\theta})\hat{A}_{t},\text{clip}\left(\bm{r}_{t}^{i}(\bm{\theta}),1\pm\epsilon\right)\hat{A}_{t}\right],$ (20)
$\text{where}\ \bm{r}_{t}^{i}(\bm{\theta})=\frac{\pi_{\theta^{i}}^{i}\left(a_{t}^{i}|\hat{o}_{t}^{i},\hat{\bm{a}}_{t}^{\mathcal{N}(i)}\right)}{\pi_{\theta^{i}_{\text{old}}}^{i}\left(a_{t}^{i}|\hat{o}_{t}^{i},\hat{\bm{a}}_{t}^{\mathcal{N}(i)}\right)},\;\mathcal{N}(i)=\left\{j\neq i\mid\mathcal{D}(a^{j},a^{i})=1\right\}.$
Action Normalization
@@
As shown in equation (21), the probability of each token $P\left(w|\cdot\right)$ is less than (or equal to) 1. As a result, longer action tends to have a lower joint probability, even though it is more reasonable than other actions in the task. To address the issue, Tan et al. (2024) proposed token normalization and word normalization by dividing the joint action probability in (21) by a constant related to its token or word length. It gets rid of the strong assumption that all the actions have similar lengths, which in LLMs, can be an evil condition, because in agentic tasks, it is all too common that solutions generated can be either super long with a chain-of-thought or super short with a direct answer.
$\text{log}P_{\pi^{i}}\left(a^{i}|\hat{o}^{i},\bm{a}^{\mathcal{N}(i)}\right)$ $=\text{log}P_{\pi^{i}}\left(w^{i,1},\ldots,w^{i,L^{i}}|\hat{o}^{i},\bm{a}^{\mathcal{N}(i)}\right)$
$=\sum_{j=1}^{L^{i}}\text{log}P_{\pi^{i}}\left(w^{i,j}|\hat{o}^{i},\bm{a}^{\mathcal{N}(i)},w^{i,1},\ldots,w^{i,j-1}\right).$ (21)
Agent-by-agent Update
@@
Updating every policy during each training epoch may introduce implicit off-policy issues, which can potentially compromise the monotonic improvement guarantee of PPO for individual agents’ policies. This problem can be exacerbated in LaMAS, where agents often have dependencies on the results of other agents’ actions. Drawing on the approach of A2PO (Wang et al., 2023), an agent-by-agent update scheme can be employed to restore monotonic improvement for each agent and facilitate stable convergence.
Trajectory-level Optimization
@@
For language models, the definition of action from the RL perspective can be dynamic. Direct trajectory-level optimization, such as GRPO with outcome supervision (Shao et al., 2024) and its variant DAPO (Yu et al., 2025), can be viewed as an action-level optimization with non-zero reward only at the end of the trajectory in the MDP or a normal action-level optimization in a single-step variant of the MDP. Consequently, trajectory-level optimization can be seamlessly integrated into MARFT by following the same idea.
4.1.2 Token-level Fine-Tuning
Naive Token-level MARFT
@@ 4.1.2 Token-level Fine-Tuning
On top of the credit assignment of each agent on its action, we can also give a more fine-grained credit assignment by naturally treating every token as an action. In this situation, we need to define the token-level multi-agent observation-action value function and observation value function as
$Q^{\bm{\pi}}\left(\bm{o}_{t},\bm{a}_{t}^{i-1},w_{t}^{i,j-1},w_{t}^{i,j}\right)\coloneqq\mathbb{E}_{\tau\sim p_{\bm{\pi}}\left(\tau|\bm{o}_{t},\bm{a}_{t}^{i-1},w_{t}^{i,j-1},w_{t}^{i,j}\right)}\left[\sum_{k=t}^{\infty}\gamma^{k-t}R_{k}(\bm{o}_{k},\bm{a}_{k})|\bm{o}_{t},\bm{a}_{t}^{i-1},w_{t}^{i,j-1},w_{t}^{i,j}\right],$ (22)
$V^{\bm{\pi}}\left(\bm{o}_{t},\bm{a}_{t}^{i-1},w_{t}^{i,j-1}\right)\coloneqq\mathbb{E}_{\tau\sim p_{\bm{\pi}}\left(\tau|\bm{o}_{t},\bm{a}_{t}^{i-1},w_{t}^{i,j-1}\right)}\left[\sum_{k=t}^{\infty}\gamma^{k-t}R_{k}(\bm{o}_{k},\bm{a}_{k})|\bm{o}_{t},\bm{a}_{t}^{i-1},w_{t}^{i,j-1}\right].$ (23)
And the token-level Bellman backup can be spontaneously derived as
$Q^{\bm{\pi}}(\bm{o}_{t},\bm{a}_{t}^{i-1},w_{t}^{i,1:j-1},w_{t}^{i,j})\leftarrow\begin{cases}0+\gamma\max\limits_{w_{t}^{i,j+1}}Q^{\bm{\pi}}(\bm{o}_{t},\bm{a}_{t}^{i-1},w_{t}^{i,1:j},w_{t}^{i,j+1})&\mathrm{if~{}}j<|a_{t}^{i}|\\ 0+\gamma\max\limits_{w_{t}^{i+1,1}}Q^{\bm{\pi}}(\bm{o}_{t},\bm{a}_{t}^{1:i},w_{t}^{i+1,1})&\mathrm{if~{}}j=|a_{t}^{i}|~{}\&~{}{i+1<n}\\ R(\bm{o}_{t},\bm{a}_{t}^{1:i})+\gamma\max\limits_{w_{t+1}^{1,1}}Q^{\bm{\pi}}(\bm{o}_{t+1},\bm{a}_{t}^{1:i},w_{t+1}^{1,1})&\mathrm{if~{}}j=|a_{t}^{i}|~{}\&~{}i=n\end{cases},$ (24)
$V^{\bm{\pi}}(\bm{o}_{t},\bm{a}_{t}^{i-1},w_{t}^{i,1:j})\leftarrow\begin{cases}0+\gamma V^{\bm{\pi}}(\bm{o}_{t},\bm{a}_{t}^{1:i-1},w_{t}^{i,1:j+1})&\mathrm{if~{}}j<|a_{t}^{i}|\\ 0+\gamma V^{\bm{\pi}}(\bm{o}_{t},\bm{a}_{t}^{1:i},w_{t}^{i+1,1})&\mathrm{if~{}}j=|a_{t}^{i}|~{}\&~{}{i<n}\\ R(\bm{o}_{t},\bm{a}_{t}^{1:i})+\gamma V^{\bm{\pi}}(\bm{o}_{t+1},\bm{a}_{t}^{1:i},\emptyset)&\mathrm{if~{}}j=|a_{t}^{i}|~{}\&~{}i=n\end{cases}.$ (25)
@@
This kind of modified token-level Bellman backup guides the computations of TD errors and the advantages of tokens, thus leading to a token-level value training and policy optimization. In the context of the naive token-level MARFT approach, each token is treated as an individual action, thereby introducing a new POMDP formulation, where the reward function is highly sparse and assigns zero rewards to all intermediate tokens and preceding agents, with the reward being allocated solely to the final token of the entire multi-agent system. Although this sparsity in the reward signal poses challenges for learning the value function, it significantly reduces the optimization complexity from $O\left(\lvert\mathcal{V}\rvert^{L}\right)$ to $O\left(\lvert\mathcal{V}\rvert\times L\right)$, as previously discussed in Section 3.3. However, due to the introduction of the different MDP, this naive token-level MARFT optimizes inconsistently with the original optimization problem (Wen et al., 2024b).

### v4

@@ Abstract
Large Language Model (LLM)-based Multi-Agent Systems (LaMAS) have demonstrated remarkable capabilities in addressing complex, agentic tasks requiring multifaceted reasoning and collaboration, from generating high-quality presentation slides to conducting sophisticated scientific research. Meanwhile, Reinforcement Learning (RL) has been widely recognized for its effectiveness in enhancing agent intelligence, but limited research has investigated the fine-tuning of LaMAS using foundational RL techniques. Moreover, the direct application of conventional Multi-Agent Reinforcement Learning (MARL) methodologies to LaMAS introduces significant challenges, stemming from the unique characteristics and mechanisms inherent to LaMAS. To address these challenges, this article presents a comprehensive study of LLM-based MARL and proposes a novel paradigm termed Multi-Agent Reinforcement Fine-Tuning (MARFT). We introduce a brand-new MG called Flex-MG, which aligns with the LaMAS optimization in real-world applications and a universal algorithmic framework tailored specifically for LaMAS, outlining the conceptual foundations, key distinctions, and practical implementation strategies. We begin by reviewing the evolution from traditional RL to Reinforcement Fine-Tuning (RFT), setting the stage for a parallel analysis in the multi-agent domain. In the context of LaMAS, we elucidate critical differences between classical MARL and MARFT, such as asynchronous agent interactions, profile-aware agent design, and heterogeneous architectural configurations etc. These differences motivate a transition toward a novel, LaMAS-oriented formulation of RFT. Central to this work is the presentation of a robust and scalable MARFT framework. We detail the core algorithm, emphasizing its modularity and adaptability, and provide a complete, open-source implementation to facilitate adoption and further research. The latter sections of the paper explore real-world application perspectives and opening challenges in MARFT, including dynamic environment modeling, sample inefficiency, and the current lack of cohesive frameworks. By bridging theoretical underpinnings with practical methodologies, this work aims to serve as a roadmap for researchers seeking to advance MARFT toward resilient, adaptive, and human-aligned solutions in agentic systems. Our implementation of the proposed framework is publicly available at: https://github.com/jwliao-ai/MARFT.
@@ 8    $\mathcal{D}\leftarrow\mathcal{D}\cup\{\left(s,a,s^{\prime},r\right)\}$
11  sample a batch $\mathcal{B}=\{(s,a,s^{\prime},r)\}\sim\mathcal{D}$
@@ 2.2 LLM-based Multi-Agent Systems (LaMAS)
  • Parameter fine-tuning. A more sophisticated approach involves updating agent parameters or network configurations to optimize performance. This paradigm integrates deep learning and specific methodologies, such as multi-agent debating to generate high-quality training datasets (Subramaniam et al., 2025) or fine-tuning programming modules for improved workflow organization (Khattab et al., 2024). CORY (Ma et al., 2024) duplicates the LLM into two agents, i.e., a pioneer and an observer, that interact through knowledge transfer and role exchange, improving training stability and performance. However, many parameter fine-tuning methods lack a theoretical foundation in RL, relying instead on ad-hoc techniques. SiriuS (Zhao et al., 2025a), a self-improving, multi-agent optimization framework using bootstrapped reasoning, iteratively refines agent behaviors by leveraging successful reasoning trajectories and augmenting unsuccessful ones. The most related work, MAPoRL (Park et al., 2025), rooted in MARL, pioneers in applying multi-agent PPO to post-train LLMs, but does not consider the unique characteristics of LaMAS and assumes homogeneous agent behavior through collaborative debate, diverging from real-world scenarios where agents often perform diverse or orthogonal roles to solve complex problems. This article will focus on RFT and MARL-related methodologies in subsequent sections, providing a principled paradigm for general LaMAS reinforcement.
@@ 2.3.2 Main Approaches of MARL
(Multi-Agent Advantage Decomposition Theorem (Kuba et al., 2021)). For any predefined permutation of $n$ agents, for any state $s\in\mathcal{S}$ and joint action $\bm{a}\in\bm{\mathcal{A}}$, the following always holds:
$A^{\bm{\pi}}(s,\bm{a}^{1:n})=\sum_{m=1}^{n}A^{\bm{\pi}}(s,\bm{a}^{1:m-1},\bm{a}^{m}).$ (17)
@@
$\hskip-5.0ptL(\theta)=\frac{1}{nT}\sum_{i=1}^{n}\sum_{t=0}^{T-1}\min\left[\bm{r}_{t}^{i}(\theta)\hat{A}_{t},\text{clip}\left(\bm{r}_{t}^{i}(\theta),1\pm\epsilon\right)\hat{A}_{t}\right],\text{where}\ \bm{r}_{t}^{i}(\theta)=\frac{\pi_{\theta}^{i}\left(a_{t}^{i}|\hat{\bm{o}}_{t}^{1:n},\hat{\bm{a}}_{t}^{1:i-1}\right)}{\pi_{\theta_{\text{old}}}^{i}\left(a_{t}^{i}|\hat{\bm{o}}_{t}^{1:n},\hat{\bm{a}}_{t}^{1:i-1}\right)}.$ (19)
@@ 3 MARFT Overview
3.1 Problem Statement and Flex-MG
@@
Figure 2: A detailed illustration of the dynamics of a Flex-MG. The dependency function (dashed purple line) can vary across timesteps.
To enhance problem formulation and accommodate MARFT, we propose Flexible Markov Game (Flex-MG), illustrated in Figure 2 and denoted as $\langle\mathcal{V},\mathcal{N},{\mathcal{S}},\bm{\mathcal{A}},\mathcal{T},\mathcal{R},\gamma,\mathcal{D}\rangle$. Here, $\mathcal{V}$ is the vocabulary, $\mathcal{N}=\{1,...,n\}$ is the agent composition of a LaMAS with some organization, i.e., some order constraint among agents, $\mathcal{S}$ is the state space, and $\bm{\mathcal{A}}=\prod^{n}_{i=1}\mathcal{A}^{i}$ is the joint action space. The transition function $\mathcal{T}\colon\mathcal{S}\times\bm{\mathcal{A}}\rightarrow\mathcal{S}^{\prime}$, written as $P(s^{\prime}|s,\bm{a})$, models state changes given joint actions. The reward function $\mathcal{R}\colon\mathcal{S}\times\bm{\mathcal{A}}\rightarrow\mathbb{R}$, denoted $R(s,\bm{a})$, assigns rewards based on states and joint actions. The discount factor $\gamma$ accounts for temporal reward decay.
Flex-MG is distinguished by its introduction of a dependency function $\mathcal{D}\colon\bm{\mathcal{A}}\times\bm{\mathcal{A}}\rightarrow\{0,1\}$, which explicitly models inter-agent dependencies. Specifically, $\mathcal{D}(a^{i},a^{j})=1$ indicates that the action of agent $j$ is conditionally dependent on the action of agent $i$. As a result, the decision-making process of agent $j$ is influenced not only by the global state $s$ but also by the actions of all agents $k$ such that $k\in\{l|\mathcal{D}(a^{l},a^{j})=1\}$. These dependencies can be integrated into agent $j$’s input via concatenation or other fusion methods, thereby forming an enriched input for the agent’s policy. Crucially, the dependency function $\mathcal{D}$ can vary across timesteps, potentially governed by an orchestration agent or a dynamic routing mechanism. This flexibility captures the evolving coordination structures in LaMAS and embodies the "flexible" nature of the proposed framework. Notably, when $\mathcal{D}(a^{i},a^{j})=0$ for all $i,j$, Flex-MG reduces to a standard multi-agent decentralized setting where agents act independently.
@@ 3.3 Differences and Resulting Challenges
Execution Asynchronicity Synchronous Asynchronous
Utilities Task-specific only Agentic with original language ability
@@
Heterogeneity Parameters In-out formats, externals, parameters
System Organization Static Dynamic
@@
Execution Asynchronicity
One of the most significant differences between conventional MARL and MARFT is the nature of agent action executions. In traditional cooperative MARL, agents’ actions are typically executed simultaneously. However, in MARFT, they are executed asynchronously. In some cases, the actions of certain agents may even depend on the outcomes of other agents’ actions. For example, in a collaborative coding assistant system, one LLM agent generates a function prototype, and another agent asynchronously refines it based on the first agent’s output, demonstrating both asynchronicity and result dependency. As a result, conventional MARL methods, which assume synchronous actions, may not be directly applicable or effective in LLM-based multi-agent systems.
@@ 4 Multi-Agent Reinforcement Fine-Tuning
4.1 MARFT Instantiations
@@
Inheriting the SM-style re-modeling via the multi-agent advantage decomposition theorem introduced in the last section, MARFT follows the procedure demonstrated in Figure 3. Given any organizational or task-solving workflow of a LaMAS, theorem (17) allows MARFT to reframe it as a sequential decision-making process. When collecting interactive trajectories, the LaMAS generates actions with an auto-regressive problem-solving style. The "encoder" constructs local roll-out states using agent profiles and other relevant information, while the "decoder" combines these states with dependent predecessors’ actions to generate the current agent’s action $a^{m}\sim p_{\pi^{m}}(a^{m}|s,a^{1:m-1})$.
Figure 3: The procedure of Multi-Agent Reinforcement Fine-Tuning. Inference and training are conducted in an alternating manner. Each agent within the MAS can have its own private APIs, tool pool, database, and other resources.
MARFT-A
To instantiate the framework, we further present Action-level MARFT (MARFT-A), a variant designed for targeted, action-level policy optimization in LaMAS. During training, MARFT-A adopts an optimization objective in (20) [^4] Our current MARFT instantiation employs a sequential decision-making process for scalability. Extending it to handle the parallel rollouts permitted by Flex-MG (i.e., a general DAG structure) would necessitate a more complex optimization objective to tackle off-policy problems, which we leave as a direction for future work.. Notably, in MARFT-A, each agent optimizes within a trust region conditioned on predecessor actions, akin to MAT, thus ensuring monotonic improvement as in HAPPO’s sequential update scheme.
$L(\bm{\theta})=\frac{1}{nT}\sum_{i=1}^{n}\sum_{t=0}^{T-1}\min\left[\bm{r}_{t}^{m}(\bm{\theta})\hat{A}_{t},\text{clip}\left(\bm{r}_{t}^{m}(\bm{\theta}),1\pm\epsilon\right)\hat{A}_{t}\right],$ (20)
$\text{where}\ \bm{r}_{t}^{m}(\bm{\theta})=\frac{\pi_{\theta^{m}}^{m}\left(a_{t}^{m}|s_{t},\hat{\bm{a}}_{t}^{1:m-1}\right)}{\pi_{\theta^{m}_{\text{old}}}^{m}\left(a_{t}^{m}|s_{t},\hat{\bm{a}}_{t}^{1:m-1}\right)}.$
@@
Action Normalization. As shown in equation (21), the probability of each token $P\left(w|\cdot\right)$ is less than (or equal to) 1. As a result, longer action tends to have a lower joint probability, even though it is more reasonable than other actions in the task. To address the issue, Tan et al. (2024) proposed token normalization and word normalization by dividing the joint action probability in (21) by a constant related to its token or word length. It gets rid of the strong assumption that all the actions have similar lengths, which in LLMs, can be an evil condition, because in agentic tasks, it is all too common that solutions generated can be either super long with a chain-of-thought or super short with a direct answer.
$\text{log}P_{\pi^{i}}\left(a^{i}|s,\bm{a}^{\mathcal{N}(i)}\right)$ $=\text{log}P_{\pi^{i}}\left(w^{i,1},\ldots,w^{i,L^{i}}|s,\bm{a}^{\mathcal{N}(i)}\right)$
$=\sum_{j=1}^{L^{i}}\text{log}P_{\pi^{i}}\left(w^{i,j}|s,\bm{a}^{\mathcal{N}(i)},w^{i,1},\ldots,w^{i,j-1}\right).$ (21)
@@
Agent-by-agent Update. Updating every policy during each training epoch may introduce implicit off-policy issues, which can potentially compromise the monotonic improvement guarantee of PPO for individual agents’ policies. This problem can be exacerbated in LaMAS, where agents often have dependencies on the results of other agents’ actions. Drawing on the approach of A2PO (Wang et al., 2023), an agent-by-agent update scheme can be employed to restore monotonic improvement for each agent and facilitate stable convergence.
@@
Trajectory-level Optimization. For language models, the definition of action from the RL perspective can be dynamic. Direct trajectory-level optimization, such as GRPO with outcome supervision (Shao et al., 2024) and its variant DAPO (Yu et al., 2025), can be viewed as an action-level optimization with non-zero reward only at the end of the trajectory in the MDP or a normal action-level optimization in a single-step variant of the MDP. Consequently, trajectory-level optimization can be seamlessly integrated into MARFT by following the same idea.
MARFT-T
@@
On top of the credit assignment of each agent on its action, we can also give a more fine-grained credit assignment by naturally treating every token as an action, leading to another token-level instantiation of MARFT named MARFT-T. In this situation, we need to define the token-level multi-agent state value functions as
$Q^{\bm{\pi}}\left(s_{t},\bm{a}_{t}^{i-1},w_{t}^{i,j-1},w_{t}^{i,j}\right)\coloneqq\mathbb{E}_{\tau\sim p_{\bm{\pi}}\left(\tau|s_{t},\bm{a}_{t}^{i-1},w_{t}^{i,j-1},w_{t}^{i,j}\right)}\left[\sum_{k=t}^{\infty}\gamma^{k-t}R_{k}(s_{k},\bm{a}_{k})|s_{t},\bm{a}_{t}^{i-1},w_{t}^{i,j-1},w_{t}^{i,j}\right].$
$V^{\bm{\pi}}\left(s_{t},\bm{a}_{t}^{i-1},w_{t}^{i,j-1}\right)\coloneqq\mathbb{E}_{\tau\sim p_{\bm{\pi}}\left(\tau|s_{t},\bm{a}_{t}^{i-1},w_{t}^{i,j-1}\right)}\left[\sum_{k=t}^{\infty}\gamma^{k-t}R_{k}(s_{k},\bm{a}_{k})|s_{t},\bm{a}_{t}^{i-1},w_{t}^{i,j-1}\right].$
And then the token-level Bellman backup can be spontaneously derived as
$Q^{\bm{\pi}}(s_{t},\bm{a}_{t}^{i-1},w_{t}^{i,1:j-1},w_{t}^{i,j})\leftarrow\begin{cases}0+\gamma\max\limits_{w_{t}^{i,j+1}}Q^{\bm{\pi}}(s_{t},\bm{a}_{t}^{i-1},w_{t}^{i,1:j},w_{t}^{i,j+1})&\mathrm{if~}j<|a_{t}^{i}|\\ 0+\gamma\max\limits_{w_{t}^{i+1,1}}Q^{\bm{\pi}}(s_{t},\bm{a}_{t}^{1:i},w_{t}^{i+1,1})&\mathrm{if~}j=|a_{t}^{i}|~\&~{i+1<n}\\ R(s_{t},\bm{a}_{t}^{1:i})+\gamma\max\limits_{w_{t+1}^{1,1}}Q^{\bm{\pi}}(s_{t+1},\bm{a}_{t}^{1:i},w_{t+1}^{1,1})&\mathrm{if~}j=|a_{t}^{i}|~\&~i=n\end{cases},$
$V^{\bm{\pi}}(s_{t},\bm{a}_{t}^{i-1},w_{t}^{i,1:j})\leftarrow\begin{cases}0+\gamma V^{\bm{\pi}}(s_{t},\bm{a}_{t}^{1:i-1},w_{t}^{i,1:j+1})&\mathrm{if~}j<|a_{t}^{i}|\\ 0+\gamma V^{\bm{\pi}}(s_{t},\bm{a}_{t}^{1:i},w_{t}^{i+1,1})&\mathrm{if~}j=|a_{t}^{i}|~\&~{i<n}\\ R(s_{t},\bm{a}_{t}^{1:i})+\gamma V^{\bm{\pi}}(s_{t+1},\bm{a}_{t}^{1:i},\emptyset)&\mathrm{if~}j=|a_{t}^{i}|~\&~i=n\end{cases}.$
@@
This type of modified token-level Bellman backup guides the computations of TD errors and the advantages of tokens, thereby leading to token-level value training and policy optimization. In the context of the MARFT-T approach, each token is treated as an individual action, thereby introducing a new MG formulation, where the reward function is highly sparse and assigns zero rewards to all intermediate tokens and preceding agents, with the reward being allocated solely to the final token of the entire multi-agent system. Although this sparsity in the reward signal poses challenges for learning the value function, it significantly reduces the optimization complexity from $O\left(\lvert\mathcal{V}\rvert^{L}\right)$ to $O\left(\lvert\mathcal{V}\rvert\times L\right)$, which is discussed in Section 3.3. However, due to the introduction of the different MG, MARFT-T optimizes inconsistently with the original optimization problem (Wen et al., 2024b).